how many feet of fencing is needed for a pool 25 ft by 20 ft with 4 ft from poolside?

Answer by jorel555(1290)

You need 4 ft. of fencing all around the pool, so you have to add 8 ft to each dimension(4 ft. on each side. Your fencing is then 33x28. Your perimeter is 2*33+2*28=66+56=122 ft of fencing needed to surround your pool!!!!!!!!!!